### v1

@@ 4.3.3 Reduction
An alternative approach to addressing optimization problems with specific objectives involves reducing them to other problems and employing specialized solvers. A prevalent method utilizes phase polynomials to represent quantum circuits, subsequently converting these polynomials into Reed-Muller codes. This transformation effectively reduces the circuit optimization problem to a Reed-Muller decoding problem, enabling the application of dedicated solving techniques. In optimizing for T gates, phase polynomials are frequently utilized to characterize circuits. In a phase function, as shown in Equation 6, introducing a common factor of $\frac{\pi}{4}$ ensures that each coefficient $\theta_{i}\in\mathbb{Z}_{8}$, where the coefficients represent the rotation parameters of Rz gates in the range $[0,2\pi]$. Even coefficients indicate that the current phase can be realized using S gates, meaning the number of T gates directly correlates with the number of odd coefficients. When these coefficients are further reduced modulo 2, where 1 and 0 represent odd and even coefficients, respectively, we seek a set $\delta\bm{\theta}\in\{0,1\}^{2^{n}-1}$ such that $\bm{\theta}\oplus\delta\bm{\theta}$ is equivalent to $\bm{\theta}$ in the circuit (where $\oplus$ denotes bitwise XOR), and the Hamming weight of the new circuit is minimized, i.e., the number of 1s is minimized. This problem is equivalent to finding a binary code in an equivalence class that has the smallest Hamming distance from $\bm{\theta}$. Amy et al. [9] demonstrate that for circuits containing only T and CNOT gates, the required code size is $\mathcal{RM}(n-4,n)^{*}$, providing comprehensive theoretical proofs for related complexity bounds. Heyfron et al. [75] propose a compilation optimization paradigm for Clifford+T sets, specifically targeting T gate optimization. Initially, non-Clifford phases are extracted by introducing redundant qubits, isolating the Clifford parts in the circuit. The optimization of non-Clifford phases is reduced to RM decoding and low-rank decomposition problems. Prior to reducing to RM code decoding, Amy et al. [7] also attempt to reduce the optimization problem for T gates to quasi-matrix partitioning problems, using the same problem setup as [9] but constructing the phase function set as a finite quasi-matrix, thereby reducing the problem of generating optimal T depth circuits to finding minimal quasi-matrix partitions.
@@
Another commonly used method is to reduce the circuit optimization problem to a SAT problem for solving. Schneider et al. [166] utilized the conclusion from [1] that an optimal Clifford circuit requires at most $\mathcal{O}(\frac{n^{2}}{\log n})$ quantum gates, and mainly focused on the synthesis and optimization of Clifford circuits. Assuming the optimal circuit contains $m$ gates, where $0<m\leq c\frac{n^{2}}{\log n}$ and $c$ is a constant, at each step $k$, $0<k\leq m$, there are $1+2n+2\times\frac{n(n-1)}{2}=1+n+n^{2}$, i.e., $\mathcal{O}(n^{2})$ ways to add quantum gates to an existing circuit. This implies a total of $m(1+n+n^{2})$, i.e., $\mathcal{O}(\frac{n^{4}}{\log n})$ binary variables are needed to represent the composition of an optimal Clifford circuit, thereby transforming the circuit optimization problem into a SAT problem for resolution. To ensure the optimal $m$, a binary search strategy can be employed to iterate through the possible values of $m$, given that $m$ ranges from $1$ to $\mathcal{O}(\frac{n^{2}}{\log n})$. Initially, set $m$ to the midpoint of this range; if a SAT solver returns a result, set the upper bound to $m$; otherwise, set the lower bound to $m+1$. This method effectively determines the optimal $m$. Moreover, the algorithm can also enhance SAT problem constraints for CNOT gate optimization.
4.3.4 Machine Learning
@@
In complement to traditional algorithms, machine learning and associated techniques have been incorporated into quantum logic circuit optimization. Although machine learning algorithms may not provide theoretical guarantees on solution accuracy, they demonstrate capability in addressing larger-scale problems, offering feasible solutions at reduced computational costs. This attribute imparts practical value to quantum logic circuit optimization algorithms, expanding their applicability in more complex scenarios.
@@
Heuristic Algorithms: The fundamental principle of heuristic algorithms is to search the solution space through a series of rules and heuristic strategies to find near-optimal solutions. Compared to traditional exact algorithms, heuristic algorithms are generally more efficient, especially when dealing with quantum logic circuit optimization problems with large solution spaces, offering near-optimal solutions at lower computational costs. Amy et al. [6] proposes a relatively primitive heuristic algorithm to optimize CNOT counts in phase polynomial representations. It constructs a ${0,1}$ matrix based on XOR strings in the phase function (see equation  6), where each column represents an XOR string from the phase function, with matrix elements set to 1 for corresponding qubits in the XOR string. Adding CNOT gates causes XOR flips in the control qubit row based on the target qubit row. At each step, the algorithm greedily selects the qubit pair that causes the most flips to add a CNOT gate. If a column has only two $1$, it deterministically adds that CNOT gate. This constructs a pure CNOT circuit and analyzes the optimal CNOT count solution. Davis et al. [42] employs the $A^{*}$ algorithm to optimize CNOT counts in circuits. Given a set of quantum gates, allowed precision error, and maximum CNOT count, the algorithm starts from an empty circuit. Candidate circuits are stored in a priority queue. In each iteration, the algorithm pops the top circuit, attempts to append gates from the given set to the circuit’s end, calculates the distance to the target unitary, and checks if it’s within the allowed error range or if the CNOT count exceeds the limit. If requirements are met, it returns the current circuit and terminates. If the queue is empty, no solution exists. The comparison function of the priority queue is the same as the total cost function of $A^{*}$ algorithm, comprising the cost of the existing path and a heuristic estimate of the additional cost needed to reach the target unitary.
@@
Reinforcement Learning: In quantum logic circuit optimization, reinforcement learning offers a significant advantage over other algorithms by automatically discovering novel or undefined optimization strategies and matchable patterns. Fosel et al. [56] present a preliminary algorithm that encodes quantum gates as triplets within a three-dimensional tensor representation of quantum circuits, using qubit indices, gate types, and depth as coordinates. The algorithm incorporates predefined hard rules (transformations with direct benefits, such as adjacent gate cancellation) and soft rules (transformations with indirect potential benefits, like swapping adjacent commutable gates). During each iteration, the agent is limited to deciding on soft rule usage, while hard rules are applied to the maximum extent for optimal benefit. Agent actions are tensor-encoded, with single-gate rules encoded by depth and qubit index and two-gate rules by the minimum common qubit index and the depth of the first gate. The reinforcement learning-trained agent employs a convolutional neural network, taking the circuit’s 3D tensor representation as the input state and outputting the policy (encoded circuit transformation actions) and the current state value (indicating optimization potential). Rewards are based on cost changes, such as the reduction in the CNOT count, before and after each transformation.
@@
Li et al. [110] replace manually specified action sets with a policy network. It first converts gate circuits into graph structures and employs a graph neural network to extract quantum gate representations. The policy network includes quantum gate and transformation selection, incorporating probability distribution-based sampling with gradient updates via temperature-coefficient softmax. Experimental results show significant improvements over existing compilation optimizers. Riu et al. [161] introduce a reinforcement learning method for ZX calculus, using a policy network in combination with a graph neural network to identify replaceable positions in ZX diagrams. Actions involve substituting the output positions from the policy network, where the substitution concentrates on common equivalent circuit simplifications for the Clifford set. Rewards are based on gate count, necessitating the conversion of ZX diagrams to quantum circuits. Delayed conversion strategies are employed to reduce computational overhead. Ruiz et al. [163] optimize T count by transforming the problem into the low-rank decomposition of three-dimensional tensors, adopting Google’s AlphaTensor method [53]. It employs phase polynomials up to the third order, corresponding to T gates, Controlled-S gates, and Controlled-Controlled-Z gates. By encoding these coefficients into a third-order tensor, the problem is reframed as the optimal decomposition of the circuit’s third-order tensor. Additional pattern matching strategies for T and Toffoli gates are employed to streamline consecutive circuit segments. The study highlights that reinforcement learning algorithms can spontaneously discover novel matching patterns, complementing the manually designed strategies.
4.3.5 Circuit Optimization for Variational Quantum Algorithms
@@
Variational Quantum Algorithms (VQAs) are considered among the most promising candidates for demonstrating quantum supremacy in the NISQ era, owing to their shallow circuit depth and the adaptability to circuit noise and other factors. Consequently, numerous quantum circuit optimization algorithms have been developed, targeting specific VQAs and corresponding quantum hardware characteristics. Several comprehensive reviews have explored related issues [180, 54, 31, 24]. This section presents a concise overview of these recent developments in the field.
@@
Variational Quantum Eigensolver (VQE): A key application of VQE is solving ground state energies in quantum chemistry problems. Common algorithms include Hardware-Efficient Ansatz (HEA) [90] and Unitary Coupled Cluster (UCC) [162]. HEA circuits adhere to hardware constraints and are simple but less accurate, while UCC methods are precise but complex due to creation-annihilation operator implementations. Continuous efforts have been focused on simplifying these operators in circuits. ADAPT-VQE [63] reduces operator usage by incrementally building ansatz circuits, selecting operators with the highest energy gradient from a pre-determined pool. This accelerates VQE convergence and significantly reduces the number of required operators and circuit depth. Qubit-ADAPT-VQE [177] implements a hardware-friendly ADAPT-VQE variant to minimize SWAP gates in existing quantum hardware topologies. QEB-ADAPT-VQE [204] uses Qubit-excitation evolution instead of Fermionic-excitation, allowing fixed structures for single and double excitation operators. It outperforms FEB-based algorithms [63, 177] in CNOT count, circuit parameter count, and algorithm iterations. Magoulas et al. [117] combines FEB and QEB methods, employing an approximation strategy to balance CNOT gate reduction and accuracy preservation. DISCO-VQE [28] achieves more compact ansatz circuits by combining spin-symmetry-preserving fermionic operator products with a global optimization algorithm, simultaneously optimizing discrete operators and continuous variational parameters. TETRIS-ADAPT-VQE [10] accelerates ansatz synthesis by introducing multiple non-interfering operators per iteration, contrasting with ADAPT-VQE’s single-operator approach. Diverging from ADAPT methods, Halder et al. [68] incorporates Restricted Boltzmann Machines (RBM) [5, 106, 164] and many-body perturbation theory in ansatz construction. It uses low-level excitation UCC operators as the RBM training set to generate higher-level excitation operators, achieving high accuracy while maintaining low complexity (e.g., UCCSDT accuracy with UCCSD-level circuit complexity).
@@
Quantum Approximate Optimization Algorithm: The Quantum Approximate Optimization Algorithm (QAOA), introduced by Farhi et al. [52], plays a crucial role in solving quadratic unconstrained binary optimization (QUBO) problems. Alam et al. [3] proposes four general methods for optimizing QAOA circuits through gate commutation. Focusing specifically on QAOA circuits designed to solve the Max-Cut problem, Arufe et al. [11] presents a decomposition-based genetic algorithm for circuit optimization. Majumdar et al. [119] introduces two optimization strategies based on edge coloring [190, 186] and depth-first search, respectively, yielding optimizations of $\lfloor\frac{n}{2}\rfloor$ and $n-1$ CNOT gates (where $n$ represents the number of vertices in the original graph). Majumdar et al. [118] further improves upon the DFS-based strategy from [119], proposing a heuristic greedy algorithm that reduces circuit depth while maintaining the $n-1$ CNOT gate optimization. Inspired by the ansatz construction in ADAPT-VQE [63, 177], ADAPT-QAOA [215] selects operators with the highest energy gradient from an operator pool in each iteration, accelerating algorithm convergence. ADAPT-QAOA adapts to various hardware constraints, increasing the approximation ratio while reducing circuit parameters and CNOT count. Furthermore, Dynamic-ADAPT-QAOA [200] builds upon ADAPT-QAOA by eliminating layers with optimal parameters approaching zero, further reducing QAOA circuit size. Herrman et al. [74] approach QAOA circuit optimization from the perspective of enhancing combinatorial optimization problems, providing a global optimization method.
5 Qubit Mapping and Routing
@@
To execute quantum circuits on quantum computers, logical qubits in the circuit must be mapped to the physical qubits of the quantum processor. Since physical qubits are not fully connected, two-qubit (and multi-qubit) gates in the circuit may not be directly deployable after mapping, necessitating the introduction of additional SWAP gates to satisfy connectivity constraints. SWAP gates exchange information and states between two qubits, enabling two-qubit gates between non-adjacent qubits. Implementing a SWAP gate requires three CNOT gates. In the NISQ era, CNOT gates still exhibit significant noise, and the excessive introduction of SWAP gates can severely affect execution results. Therefore, minimizing the number of SWAP gates introduced is crucial. Figure 3 illustrates an example where, based on the given superconducting quantum topology, physical qubits $Q_{9}$ and $Q_{10}$ cannot be connected, preventing the direct application of two-qubit gates between them. When compiling the logical circuit shown in Figure 3(b), three SWAP gates must be added to exchange qubits, ensuring the circuit satisfies the connectivity in Figure 3(a) (the last SWAP is omitted as the measurement step is unaffected by SWAP gates). After optimizing qubit routing, the circuit can be simplified to use only one SWAP gate, as shown in Figure 3(d).
Problem Definition: The Qubit Mapping problem can be uniformly defined as: given a quantum circuit, the coupling graph of the quantum computer, and noise conditions, determine the mapping relationship between logical and physical qubits, such that the mapped circuit satisfies connectivity constraints while minimizing the mapping’s impact on circuit execution results.
5.1 Qubit Mapping and Routing Targets
5.1.1 Gate Count

### v2

@@ IV-C3 Reduction
Phase polynomials are often employed in optimizing T gates. In a phase function, as shown in Eq. 2, introducing a common factor of $\frac{\pi}{4}$ ensures that each coefficient $\theta_{i}\in\mathbb{Z}_{8}$, where the coefficients represent the rotation parameters of Rz gates in the range $0,2\pi]$. Even coefficients indicate that the current phase can be realized using S gates, meaning the number of T gates directly correlates with the number of odd coefficients. When these coefficients are further reduced modulo 2, where 1 and 0 represent odd and even coefficients, respectively, we seek a set $\delta\bm{\theta}\in\{0,1\}^{2^{n}-1}$ such that $\bm{\theta}\oplus\delta\bm{\theta}$ is equivalent to $\bm{\theta}$ in the circuit (where $\oplus$ denotes bitwise XOR), with the Hamming weight of the new circuit minimized. This problem is equivalent to finding a binary code in an equivalence class that has the smallest Hamming distance from $\bm{\theta}$. In [27], it is shown that for circuits containing only T and CNOT gates, the required code size is $\mathcal{RM}(n-4,n)^{*}$, providing comprehensive theoretical proofs for related complexity bounds. [26] proposed a compilation optimization paradigm for Clifford+T sets, specifically targeting T gate optimization. Non-Clifford phases were first extracted by introducing redundant qubits, isolating the Clifford parts in the circuit. The optimization of non-Clifford phases was then reduced to Reed-Muller (RM) decoding and low-rank decomposition problems. Prior to reducing to RM code decoding, [13] also attempted to reduce the optimization problem for T gates to quasi-matrix partitioning problems, using the same problem setup as [27] but constructing the phase function set as a finite quasi-matrix.
@@
Another popular method is to reduce the circuit optimization problem to a SAT problem for resolution. [28] utilized the conclusion from [201] that an optimal Clifford circuit requires at most $\mathcal{O}(\frac{n^{2}}{\log n})$ quantum gates, focusing on the synthesis and optimization of Clifford circuits. Assuming the optimal circuit contains $m$ gates, where $0<m\leq c\frac{n^{2}}{\log n}$ and $c$ is a constant, at each step $k\leq m$, there are $\mathcal{O}(n^{2})$ ways to add quantum gates to an existing circuit. This implies a total of $m(1+n+n^{2})$, i.e., $\mathcal{O}(\frac{n^{4}}{\log n})$ binary variables are needed to represent the composition of an optimal Clifford circuit, thereby transforming the circuit optimization problem into a SAT problem for resolution. To ensure the optimal $m$, a binary search strategy can be employed to iterate through the possible values of $m$, given that $m$ ranges from $1$ to $\mathcal{O}(\frac{n^{2}}{\log n})$.
IV-C4 Machine Learning
@@
In complement to traditional algorithms, machine learning and associated techniques have been incorporated into quantum logic circuit optimization. Although machine learning algorithms may not provide theoretical guarantees on solution accuracy, they demonstrate the capability in addressing larger-scale problems, offering feasible solutions at reduced computational costs. This attribute imparts practical value to quantum logic circuit optimization algorithms, expanding their applicability in more complex scenarios.
@@
Heuristic Algorithms: The principle of heuristic algorithms is to search the solution space through a series of rules and heuristic strategies to find near-optimal solutions. Compared to traditional exact algorithms, heuristic algorithms are generally more efficient, especially when dealing with quantum logic circuit optimization problems with large solution spaces, offering near-optimal solutions at lower computational costs. The work [65] proposed a relatively primitive heuristic algorithm to optimize CNOT counts in phase polynomial representations. It constructs a 0,1{0,1} matrix based on XOR strings in the phase function (see Eq. 2), where each column represents an XOR string from the phase function, with matrix elements set to 1 for corresponding qubits in the XOR string. Adding CNOT gates causes XOR flips in the control qubit row based on the target qubit row. At each step, the algorithm greedily selects the qubit pair that causes the most flips to add a CNOT gate. If a column has only two $1$, it deterministically adds that CNOT gate. This constructs a pure CNOT circuit and analyzes the optimal CNOT count solution. [165] employed the $A^{*}$ algorithm to optimize CNOT counts in circuits. Given a set of quantum gates, allowed precision error, and maximum CNOT count, the algorithm starts from an empty circuit. Candidate circuits are stored in a priority queue. In each iteration, the algorithm pops the top circuit, attempts to append gates from the given set to the circuit’s end, calculates the distance to the target unitary, and checks if it’s within the allowed error range or if the CNOT count exceeds the limit. If requirements are met, it returns the current circuit and terminates. If the queue is empty, no solution exists. The comparison function of the priority queue is the same as the total cost of $A^{*}$, comprising the cost of the existing path and a heuristic estimate of the additional cost to reach the target unitary.
@@
Reinforcement Learning: In quantum logic circuit optimization, reinforcement learning offers a significant advantage over other algorithms by automatically discovering novel or undefined optimization strategies and matchable patterns. [29] presented a preliminary algorithm that encodes quantum gates as triplets within a three-dimensional tensor representation of quantum circuits, using qubit indices, gate types, and depth as coordinates. The algorithm incorporates predefined hard rules (transformations with direct benefits, e.g., adjacent gate cancellation) and soft rules (transformations with indirect potential benefits, like swapping adjacent commutable gates). During each iteration, the agent is limited to deciding on soft rule usage, while hard rules are applied to the maximum extent for optimal benefit. Agent actions are tensor-encoded, with single-gate rules encoded by depth and qubit index and two-gate rules by the minimum common qubit index and the depth of the first gate. The reinforcement learning-trained agent employs a convolutional neural network, taking the circuit’s 3D tensor representation as the input state and outputting the policy (encoded circuit transformation actions) and the current state value (indicating optimization potential). Rewards are based on cost changes, e.g. the reduction in the CNOT count, before and after each transformation.
@@
The study in [30] replaced manually specified action sets with a policy network. It first converts gate circuits into graph structures and employs a graph neural network to extract quantum gate representations. The policy network includes quantum gate and transformation selection, incorporating probability distribution-based sampling with gradient updates via temperature-coefficient softmax. Experimental results show significant improvements over existing compilation optimizers. In [31], a reinforcement learning method is introduced for ZX calculus, using a policy network in combination with a graph neural network to identify replaceable positions in ZX diagrams. Actions involve substituting the output positions from the policy network, where the substitution concentrates on common equivalent circuit simplifications for the Clifford set. Rewards are based on gate count, necessitating the conversion of ZX diagrams to quantum circuits. Delayed conversion strategies are employed to reduce computational overhead.  [67] optimized T count by transforming the problem into the low-rank decomposition of three-dimensional tensors, adopting Google’s AlphaTensor method [202]. It employed phase polynomials up to the third order, corresponding to T gates, Controlled-S gates, and Controlled-Controlled-Z gates. By encoding these coefficients into a third-order tensor, the problem was reframed as the optimal decomposition of the circuit’s third-order tensor. Additional pattern matching strategies for T and Toffoli gates were employed to streamline consecutive circuit segments. The study highlights that reinforcement learning algorithms can spontaneously discover novel matching patterns, complementing the manually designed strategies.
IV-C5 Circuit Optimization for VQAs
@@
VQAs are considered among the most promising candidates for demonstrating quantum supremacy in the NISQ era, owing to their shallow circuit depth and the adaptability to circuit noise and other factors. Consequently, numerous quantum circuit optimization algorithms have been developed, targeting specific VQAs and corresponding quantum hardware characteristics. Several comprehensive reviews have explored related issues [203, 204, 205, 206]. This section presents a concise overview.
@@
Variational Quantum Eigensolver: A key application of VQE is solving ground state energies in quantum chemistry problems. Common algorithms include HEA [207] and Unitary Coupled Cluster (UCC) [208]. HEA circuits adhere to hardware constraints and are simple but less accurate, while UCC methods are precise but complex due to creation-annihilation operator implementations. Continuous efforts have been focused on simplifying these operators in circuits. ADAPT-VQE [18] reduced operator usage by incrementally building ansatz circuits, selecting operators with the highest energy gradient from a pre-determined pool. This accelerates VQE convergence and significantly reduces the number of required operators and circuit depth. Qubit-ADAPT-VQE [32] implemented a hardware-friendly ADAPT-VQE variant to minimize SWAP gates in existing quantum hardware topologies. QEB-ADAPT-VQE [33] used Qubit-excitation evolution instead of Fermionic-excitation, allowing fixed structures for single and double excitation operators. It outperforms FEB-based algorithms [18, 32] in CNOT count, circuit parameter count, and algorithm iterations. [34] combined FEB and QEB methods, employing an approximation strategy to balance CNOT gate reduction and accuracy preservation. DISCO-VQE [209] achieved more compact ansatz circuits by combining spin-symmetry-preserving fermionic operator products with a global optimization algorithm, simultaneously optimizing discrete operators and continuous variational parameters. TETRIS-ADAPT-VQE [193] accelerated ansatz synthesis by introducing multiple non-interfering operators per iteration, contrasting with ADAPT-VQE’s single-operator approach. Diverging from ADAPT methods, [192] incorporated Restricted Boltzmann Machines (RBM) [210] and many-body perturbation theory in ansatz construction. It utilized low-level excitation UCC operators as the RBM training set to generate higher-level excitation operators, achieving high accuracy while maintaining low complexity (e.g., UCCSDT accuracy with UCCSD-level circuit complexity).
@@
Quantum Approximate Optimization Algorithm (QAOA): As introduced by [111], it plays a crucial role in solving quadratic unconstrained binary optimization (QUBO) problems. [211] proposed four general methods for optimizing QAOA circuits through gate commutation. Focusing specifically on QAOA circuits designed to solve the Max-Cut problem, [212] presented a decomposition-based genetic algorithm for circuit optimization. The work [113] introduced two optimization strategies based on edge coloring [213, 214] and depth-first search, respectively, yielding optimizations of $\lfloor\frac{n}{2}\rfloor$ and n1n-1 CNOT gates (where $n$ represents the number of vertices in the original graph). [112] further improved upon the DFS-based strategy from [113], proposing a heuristic greedy algorithm that reduced circuit depth while maintaining the n1n-1 CNOT gate optimization. Inspired by the ansatz construction in ADAPT-VQE [18, 32], ADAPT-QAOA [194] selected operators with the highest energy gradient from an operator pool in each iteration, accelerating algorithm convergence. ADAPT-QAOA can be adapted to various hardware constraints, increasing the approximation ratio while reducing circuit parameters and CNOT count. Furthermore, Dynamic-ADAPT-QAOA [215] built upon ADAPT-QAOA by eliminating layers with optimal parameters approaching zero, further reducing QAOA circuit size. In [172], the authors approached QAOA circuit optimization from the perspective of enhancing combinatorial optimization problems, providing a global optimization method. Another aspect in optimizing the QAOA ansatz is to improve the expressivity of the single QAOA layer and thus reduce the overall layer numbers to optimize circuit depth. MA-QAOA [173] disabled the parameter sharing strategy, leading to qubit-number independent parameters for each layer. XQAOA [216] altered the mixer Hamiltonian with extra parameterized rotation gates to further enhance ansatz expressivity.
V Qubit Mapping and Routing
@@
The execution of quantum circuits on quantum processors necessitates the mapping of logical qubits to physical qubits. Given that physical qubits are not fully connected, two-qubit (and multi-qubit) gates between non-adjacent qubits cannot be directly executed after mapping. This constraint requires the insertion of additional SWAP gates to facilitate the exchange of quantum states between qubits. As a SWAP gate is composed of three CNOT gates—a primary source of noise in the NISQ era—the introduction of an excessive number of SWAP gates can significantly degrade the fidelity of circuit execution. Consequently, minimizing the SWAP gate count is a paramount objective in quantum circuit compilation. Fig. 3 provides a representative example. For the depicted superconducting quantum topology, a two-qubit gate cannot be directly applied between physical qubits $Q_{9}$ and $Q_{10}$ due to the absence of a direct link. The initial compilation of the logical circuit in Fig. 3(b) would necessitate the insertion of three SWAP gates to satisfy connectivity. Subsequent qubit routing optimization, however, can reduce the required number of SWAP gates to a single instance, as demonstrated in Fig. 3(d). The final SWAP gate is typically omitted since it does not influence the measurement outcome.
Problem Definition: The Qubit Mapping problem is defined as finding an optimal initial mapping of logical qubits to physical qubits and a corresponding qubit routing schedule. Given a quantum circuit, a quantum processor’s coupling graph, and its noise conditions, the objective is to satisfy all two-qubit gate connectivity constraints by minimizing the overhead of auxiliary SWAP gates. This minimization is crucial for maximizing the overall circuit fidelity.
V-A Qubit Mapping and Routing Targets
V-A1 Gate Count